The average number of milligrams (mg) of cholesterol in a cup of a certain brand of ice cream is 660 mg, and the standard deviation is 35 mg. Assume that the variable is normally distributed. If a sample of 10 cups of ice cream is selected, what is the probability that the mean of the sample will be larger than 670 mg?


1
Expert's answer
2021-05-11T12:01:35-0400

Solution:

Given, "\mu=660,\sigma=35,n=10"

"X\sim N(\mu,\sigma^2)"

Now, "P(X>670)=P(z>\dfrac{670-660}{35/\sqrt{10}})=P(z>0.9035)"

"=1-P(z\le0.9035)\approx1-0.81594=0.18403"